# From News to Forecast: Iterative Event Reasoning in LLM-Based Time Series Forecasting

### v1

@@ Abstract
This paper introduces a novel approach to enhance time series forecasting using Large Language Models (LLMs) and Generative Agents. With language as a medium, our method adaptively integrates various social events into forecasting models, aligning news content with time series fluctuations for enriched insights. Specifically, we utilize LLM-based agents to iteratively filter out irrelevant news and employ human-like reasoning and reflection to evaluate predictions. This enables our model to analyze complex events, such as unexpected incidents and shifts in social behavior, and continuously refine the selection logic of news and the robustness of the agent’s output. By compiling selected news with time series data, we fine-tune the LLaMa2 pre-trained model. The results demonstrate significant improvements in forecasting accuracy and suggest a potential paradigm shift in time series forecasting by effectively harnessing unstructured news data.
@@ 1 Introduction
Time series forecasting [18, 21] serves as an essential foundation for decision-making across diverse economic, infrastructural, and social domains [2, 13, 14, 16, 56]. The purpose of analyzing time series data is to decode the evolving relationships within complex and dynamic real-world systems. Traditional forecasting methods, which excel at identifying patterns in historical data, perform well when time series distributions remain consistent over time. However, they have limitations in addressing sudden disruptions or anomalies caused by external random events, and do not systematically connect complex social events with fluctuations in time series data. Incorporating insights that reflect real-world events and their impact on social and economic behaviors into time series forecasts is critical for enhancing both the reliability and accuracy of these forecasts.
News articles provide crucial insights into unexpected incidents, policy changes, technological developments, and public sentiment shifts. These are factors that numerical data alone may not reflect. Integrating news into forecasting enriches its inputs with context that closely mirrors the complexities of human behavior and societal changes. On the one hand, news offers a real-time snapshot of events, enabling the model to adjust predictions based on updated information. On the other hand, incorporating qualitative data from news sources allows the model to address non-linear and non-numeric influences, providing a comprehensive perspective on the various influences. Combining both quantitative and qualitative insights can improve forecast accuracy and reliability in rapidly changing environments. Incorporating news into forecasting not only enhances the data set but also improves the adaptability and accuracy, making the model more reflective of real-world dynamics.
In this work, we first propose a unified approach that embeds news and supplementary information into time series data using textual prompts. By fine-tuning large language models (LLMs) [54, 55, 66], we transform time series forecasting into the prediction of the next token in text, taking into account news and supplementary information. The powerful inductive reasoning capabilities of pre-trained LLMs, along with their ability to represent multi-modal distributions, enable them to perform few-shot time series predictions [17]. The potential of language models in time series forecasting has also been proven [23, 71]. After further training on our constructed dataset of time series, news, and supplementary information, these language models can produce accurate time series forecasts that consider the news and supplementary details provided in the input prompts.
@@
Effective screening of news is another key issue for enhancing time series forecasting as input diversity increases. This task requires more than simple keyword extraction; it demands a deep understanding of how news elements interact with forecast variables and extends beyond linear reasoning to include intelligent analytical methods. We employ LLM agents [43, 57] with advanced human-like reasoning capabilities for dynamic and effective news selection. These agents use few-shot learning to adapt their strategies based on insights from logical scenarios that mimic human reasoning about factors affecting time series fluctuations. This enables them to identify relevant news, which is then paired with corresponding time series data to create a context-aware dataset that improves prediction accuracy. Additionally, the LLM agent also plays a crucial role in the model evaluation process, continuously refining their selection logic by comparing forecasted outcomes with actual events. This iterative self-evaluation helps to pinpoint and integrate critical but previously overlooked news items. By automating the analysis of unstructured text and applying chain-of-thought prompting, the agents effectively uncover patterns linking news events to forecasting discrepancies, revealing the nuanced effects of external factors on predictions.
@@
  • The paper presents a novel framework capable of integrating unstructured news data with numerical time series inputs, deepening the contextual understanding of the data and enhancing the model’s ability to respond to social events and real-world dynamics.
  • We create a dataset that spans various sectors, containing not only time series data specific to certain tasks but also a curated selection of well-substantiated public news. This dataset will enable extensive exploration in future time series research.
  • The research highlights the use of LLM agents for dynamic news selection and analysis. We leverage the reasoning capabilities of LLM to automate the understanding and filtering of news content effectively. The agents continuously refine their news selection and analysis based on the forecasting results, thereby improving the models’ accuracy and reliability.
  • Our results demonstrate the effective integration of news data, achieving superior prediction accuracy across various domains such as finance, energy, traffic, and bitcoin domains. Our findings demonstrate that incorporating news is highly adept at navigating complexities, especially in energy demand patterns.
@@ 2 Related Work
Attempts to incorporate textual information (e.g., Twitter feeds, news articles, public reports) into time series forecasting have been made across various domains, such as finance [7, 48, 49], energy [3, 41], entertainment [26], pandemics [65], and tourism [47]. Traditional methods often simplify text analysis to counting keyword frequencies [41] or using dummy variables, which do not capture nuanced meanings. Advanced efforts include extracting richer textual features like word frequencies and sentiments using traditional NLP [9] and machine learning, as demonstrated by Bai et al. [3]. However, these approaches require labor-intensive feature engineering, struggle with long-text dependencies, and lack a deep contextual understanding. In contrast, large language models excel in processing complex textual data and understanding contextual relationships, which can improve prediction accuracy and efficiency through automated feature extraction and enhanced scalability across multiple tasks. Despite their potential, no study has yet fully exploited LLM for enhancing forecasting with their capabilities in understanding unstructured textual data.
Language models for time series forecasting. LLMs such as the GPT series [1, 4, 44, 45] and LLaMa [54] have excelled in a variety of natural language processing tasks. With their vast parameter sets, LLMs acquire extensive general knowledge and reasoning capabilities during pre-training, which is crucial for building intelligent systems equipped with common sense. LLMs architectures are increasingly applied to time series processing and forecasting [23, 25, 35, 37, 51]. For instance, TEMPO [6] adapts GPT architectures for dynamic temporal representation learning, while TIME-LLM [23] utilizes LLMs for time series forecasting by reprogramming input data and applying Prompt-as-Prefix techniques. Similarly, FPT [71] demonstrates that even frozen LLMs can perform effectively in time series tasks, leveraging the universality of self-attention mechanisms. Lag-LLaMa [46] uses a decoder-only transformer for univariate probabilistic forecasting, and Gruver et al. [17] show that by framing time series forecasting as next-token prediction, LLMs can surpass traditional models through effective tokenization and adaptation.
@@ 3 Method
In this work, we aim to integrate news insights into time series forecasting. The development of such a system faces several challenges. Firstly, the forecasting method must handle unstructured, non-numerical news inputs flexibly and adjust predictions based on the context of the news events. Secondly, constructing this model involves filtering news related to the time series data, establishing connections between the news and the time series data. This requires sifting through vast amounts of internet data to find relevant information, demanding deep societal understanding and sophisticated reasoning skills. Therefore, an intelligent agent is designed to manage this complexity. Moreover, potential inaccuracies in news selection or inferential errors may still affect the forecast accuracy, requiring further refinement of news selection and reasoning based on the predictions. Our approach includes three main modules: a language model-based forecasting module (Sec 3.1), a reasoning agent for news filtering and inference and an evaluation agent to assess and refine the forecasting model (Sec 3.2). The core workflow of our method and the interrelations between these modules are shown in Figure 1. The subsequent sections will discuss these three modules in detail.
@@
News context offers critical insights into complex social events that traditional data often overlook, and it also reflects sudden shifts in time series due to random events. In fact, the time series we collected can already be seen as being influenced by the aforementioned events. Assume an event $\mathcal{E}$ and a time series $x_{0:t}$, its impact on the future sequence can be expressed also as a conditional probability $P(x_{t}|x_{0:t},\mathcal{E})$. However, when information about event $\mathcal{E}$ is not provided, we can only predict through historical time series.Although time series data itself can show patterns and trends, it lacks the ability to indicate the causality behind events. The event information $\mathcal{E}$ offers the context needed to understand why certain spikes or drops occur. We show in Figure 2 some news that are closely related to time series forecasting.
In language models, news events can also be represented as text tokens in the language model. Consider a set of news text tokens $\{e_{0},e_{1},\dots,e_{u}\}$, which represent event $\mathcal{E}$. LLMs can treat this news information as a condition input and perform conditional probability predictions $P(x_{t}|x_{0:t},e_{0:u})$. Including $e_{0:u}$ provides crucial contextual information that influences the prediction of future values. This process aligns with the standard approach used by language models to interpret text, allowing for the incorporation of information about multiple events through various news contexts simultaneously to enhance prediction accuracy. In practice, we only need to integrate news text with historical time series data in a prompt engineering manner.
@@
Time series and news pre-pairing. For the initial stage of data preparation, news is retrieved to align with time series data based on matching time frequencies, horizons, and geographical areas. This synchronization ensures that insights from textual information are timely and regionally relevant. For example, to understand state-level electricity demand in Australia from 2018 to 2022, we gather local news from various Australian states and international news occurring in the same period that might directly or indirectly affect demand. In this way, potentially relevant candidate information can be roughly selected first, and such screening can be easily completed through crawler means.
@@
Reasoning agent for news selection. We employ an LLM-based reasoning agent capable of sophisticated tasks such as conversation, reasoning, and semi-autonomous action. This agent is programmed using detailed prompts that define roles, instructions, permissions, and context, enabling it to interpret human commands and perform complex tasks. This approach effectively condenses extensive news datasets into a refined selection of pertinent articles. It leverages its inferencing abilities to efficiently sift through, categorize, and interpret news texts. We employ few-shot prompting [4] combined with the CoT [57] method to develop an agent that understands the context of news for time series forecasting. This technique enhances the model’s ability to handle multi-step, commonsense reasoning tasks essential for accurate forecasting. The agent uses multi-step prompts to break down complex problems into simpler, manageable parts. Our three-phase prompting method, illustrated in Figure 3 Part I, includes: 1) we develop an understanding of time series influencers, sorting them by impact (positive/negative) and duration (short/long-term), considering economic, policy, seasonal, and technological factors; 2) we direct the agent to filter and categorize news based on either automatically generated logic or a given reasoning logic, focusing on relevance to time series and classifying the impact (e.g., long-term and short-term); 3) we specify the output format for the agent to organize the selected news into JSON, detailing aspects like summary, affected region, reporting time, and rationale. More details about our prompting method can be found in Appendix A.6. The LLM agent can automatically form the understanding of time series influencers, and providing a given reasoning logic in our models is optional. In the automated process, the agent forms its logic through prompts designed to help it determine how different types of news affect a specific domain. For example, we use open-ended questions to allow the agent to independently summarize and develop filtering logic. User knowledge can also be incorporated into these prompts as a given reasoning to help the agent generate more comprehensive logic. The agent then filters news based on the generated logic, either fully automatic or including user-provided input.
@@
Evaluation agent for reasoning updating We also design an evaluation agent to assess and improve the effectiveness of the aforementioned news filtering. Relying solely on the reasoning agent for news selection is not optimal, as the interaction between news and time series is complex. The reasoning agent can only analyze the potential impact of different news from the perspective of news content, without knowing whether the trained time series forecasting model based on them can make accurate forecasts. The evaluation agent is deployed after the time series prediction model has been trained. The evaluation agent extends beyond simple numerical assessments of prediction accuracy by incorporating human-like logical reasoning to refine the news selection logic chain. We focus our evaluation on identifying inaccuracies potentially caused by missing news, such as unusual events or illogical reports. It observes the model’s predictive outcomes to determine if any crucial news has been overlooked and adjusts the news filtering strategy for the training data based on these results.
@@
For the evaluation agent, we structured the prompt design into three phases, as illustrated in Figure 3 Part II. In the first phase, we input the forecasting task type, the time horizon, and background information, which the agent uses to generate the steps necessary for evaluating the prediction outcomes. In the second phase, we provide the ground truth and the differences between predicted series and ground truth, alongside selected and historical news data. The agent analyzes this to identify overlooked news based on the distribution of prediction errors across different periods. In the third phase, drawing on insights from these analyses, the agent generates updated logic that informs subsequent rounds of news selection. Finally, after processing all prediction data from the validation set, the reasoning agent refines all fragmented updated logic into a cohesive final logic.
@@ 3.3 Overall Pipeline
We integrate the news reasoning and evaluation agents with the fine-tuning of the LLM forecasting model to enhance the quality of training data, as illustrated in Figure 4. In the first iteration, we use the LLM agent to establish news selection logic based on the domain and timing of the time series task. This logic directs the reasoning agent to filter relevant news, align it with time series data, and input it into the model for initial fine-tuning. After validating the model’s prediction with a validation set, an evaluation agent checks for missing news that may have influenced the prediction. This feedback helps the reasoning agent refine the filtering logic in subsequent iterations. This cycle continues until the final iteration, where the reasoning agent consolidates all updates to create the definitive news filter for training the final model. We use the GPT4-Turbo model as the LLM used in the above agents.

### v2

@@ Abstract
This paper introduces a novel approach that leverages Large Language Models (LLMs) and Generative Agents to enhance time series forecasting by reasoning across both text and time series data. With language as a medium, our method adaptively integrates social events into forecasting models, aligning news content with time series fluctuations to provide richer insights. Specifically, we utilize LLM-based agents to iteratively filter out irrelevant news and employ human-like reasoning to evaluate predictions. This enables the model to analyze complex events, such as unexpected incidents and shifts in social behavior, and continuously refine the selection logic of news and the robustness of the agent’s output. By integrating selected news events with time series data, we fine-tune a pre-trained LLM to predict sequences of digits in time series. The results demonstrate significant improvements in forecasting accuracy, suggesting a potential paradigm shift in time series forecasting through the effective utilization of unstructured news data. 22footnotetext: Code and data are available at https://github.com/ameliawong1996/From_News_to_Forecast.
@@ 1 Introduction
Time series forecasting [18, 21] serves as an essential foundation for decision-making across a wide range of economic, infrastructural, and social domains [2, 13, 14, 16, 56]. The purpose of analyzing time series data is to decode the evolving relationships within complex, dynamic real-world systems. Traditional forecasting methods, while effective at identifying patterns in historical data, perform well when time series distributions remain consistent over time. However, they have limitations in addressing sudden disruptions or anomalies caused by external random events, and do not systematically connect complex social events with fluctuations in time series data. Integrating insights from real-world events and their effects on social and economic behavior is crucial for improving the reliability and accuracy of time series forecasting.
News articles provide crucial insights into unexpected incidents, policy changes, technological developments, and public sentiment shifts—factors that numerical data alone may not capture. Integrating news into forecasting enriches its inputs with context that closely mirrors the complexities of human behavior and societal changes. On the one hand, news offers a real-time snapshot of events, enabling the model to adjust predictions based on updated information. On the other hand, qualitative data from news sources enables the model to account for non-linear and non-numeric influences. By combining both quantitative and qualitative insights, the model can improve forecast accuracy, especially in rapidly changing environments, making it more reflective of real-world dynamics.
In this work, we propose a unified approach that embeds news and supplementary information into time series data using textual prompts. By fine-tuning large language models (LLMs) [54, 55, 66], we transform time series forecasting into the prediction of the next token in text, taking into account relevant contextual information. The inductive reasoning capabilities of pre-trained LLMs, along with their ability to model multi-modal distributions, enable few-shot predictions in time series [17]. The potential of language models in time series forecasting has also been proven [23, 71]. After further training on our dataset, which includes time series, news, and supplementary information, language models can generate forecasts that consider the textual context provided in the input prompts.
@@
Effective news filtering is a key issue for enhancing time series forecasting as input diversity increases. This task requires more than simple keyword extraction; it demands a deep understanding of how news elements interact with forecast variables, extending beyond linear reasoning to incorporate intelligent analytical methods. We employ LLM agents [43, 57] with advanced human-like reasoning capabilities for dynamic and effective news selection. These agents use few-shot learning to adapt their strategies based on logical scenarios that mimic human reasoning about factors affecting time series fluctuations. This enables them to identify relevant news, which is then paired with corresponding time series data to create a context-aware dataset that improves prediction accuracy. Additionally, LLM agents also play a crucial role in model evaluation, continuously refining their selection logic by comparing forecast errors with all relevant events. This iterative self-evaluation helps identify and integrate critical news items previously overlooked. By automating the analysis of unstructured text and applying chain-of-thought prompting, the agents effectively uncover patterns linking news events to forecasting discrepancies, revealing the nuanced effects of external factors on predictions.
@@
  • The paper introduces a novel time series forecasting framework that integrates unstructured news data with numerical time series inputs, providing deeper contextual understanding and improving the model’s responsiveness to social events and real-world dynamics.
  • The research highlights the use of LLM agents for dynamic news selection and analysis. We leverage the reasoning capabilities of LLM to automate the effective understanding and filtering of news content. The agents iteratively refine their news selection process based on forecasting results, improving model accuracy and reliability.
  • We propose a data construction method that integrates time series data with news information, and have built a dataset spanning multiple domains to support our research. The dataset includes task-specific time series data and verified public news reports, which facilitates further exploration in time series research.
  • Our experiment results demonstrate the effective integration of news data, achieving superior prediction accuracy across diverse domains such as finance, energy, traffic, and bitcoin domains. Our findings demonstrate that incorporating news is highly adept at navigating complexities, especially in energy demand patterns.
@@ 2 Related Work
Attempts to incorporate textual information (e.g., Twitter feeds, news articles, public reports) into time series forecasting have been made across various domains, such as finance [7, 48, 49], energy [3, 41], entertainment [26], pandemics [65], and tourism [47]. Traditional methods often simplify text analysis to counting keyword frequencies [41] or using dummy variables, which do not capture nuanced meanings. Advanced efforts include extracting richer textual features like word frequencies and sentiments using traditional NLP [9] and machine learning methods, as demonstrated by Bai et al. [3]. However, these approaches require labor-intensive feature engineering, struggle with long-text dependencies, and lack a deep contextual understanding. In contrast, LLMs excel in processing complex textual data and understanding contextual relationships, which can improve prediction accuracy and efficiency through automated feature extraction and enhanced scalability across multiple tasks. Despite their potential, no study has yet fully exploited LLM to enhance forecasting with their capabilities in understanding unstructured textual data.
Language models for time series forecasting. LLMs such as the GPT series [1, 4, 44, 45] and LLaMa [54] have excelled in a variety of natural language processing tasks. With their vast parameter sets, LLMs acquire extensive general knowledge and reasoning capabilities during pre-training, which is crucial for building intelligent systems equipped with common sense. LLMs architectures are increasingly applied to time series processing and forecasting [23, 25, 35, 37, 51]. For instance, TEMPO [6] adapts GPT architectures for dynamic temporal representation learning, while TIME-LLM [23] utilizes LLMs for time series forecasting by reprogramming input data and applying Prompt-as-Prefix techniques. Similarly, FPT [71] demonstrates that even frozen LLMs can perform effectively in time series tasks, leveraging the universality of self-attention mechanisms. Lag-LLaMa [46] uses a decoder-only transformer for univariate probabilistic forecasting, and Gruver et al. [17] show that by framing time series forecasting as next-token prediction, LLMs can surpass traditional models through effective tokenization and adaptation. However, existing studies mainly utilize the mapping capabilities of LLMs for numerical regression, without incorporating external textual inputs or leveraging the reasoning abilities of LLMs in understanding language.
@@ 3 Method
In this work, we aim to integrate news insights into time series forecasting. The development of such a system faces several challenges. Firstly, the forecasting method must handle unstructured, non-numerical news inputs flexibly and adjust predictions based on the context of the news events. Secondly, constructing this model involves filtering news and establishing connections between the news and the time series data. This requires sifting through vast amounts of internet data to find relevant information, demanding deep societal understanding and sophisticated reasoning skills. Therefore, an intelligent agent is designed to manage this complexity. Moreover, potential inaccuracies in news selection or inferential errors may still affect the forecast accuracy, requiring further refinement of news selection and reasoning based on the predictions. Our approach includes three main modules: a language model-based forecasting module (Sec 3.1), a reasoning agent for news filtering and inference, and an evaluation agent to assess and refine the forecasting model (Sec 3.2). The core workflow of our method and the interrelations between these modules are shown in Figure 1. The subsequent sections will discuss these three modules in detail.
@@
News context offers critical insights into complex social events that traditional numerical data often overlook, and it also reflects sudden shifts in time series due to random events. In fact, the time series we collected can already be seen as being influenced by the aforementioned events. Assume an event $\mathcal{E}$ and a time series $x_{0:t}$, its impact on the future sequence can be expressed also as a conditional probability $P(x_{t}|x_{0:t},\mathcal{E})$. However, when information about event $\mathcal{E}$ is not provided, we can only predict through historical time series. Although time series data itself can show patterns and trends, it lacks the ability to indicate the causality behind events. The event information $\mathcal{E}$ offers the context needed to understand why certain spikes or drops occur. We show in Figure 2 some news that are closely related to time series forecasting.
In language models, news events can also be represented as text tokens. Consider a set of news text tokens $\{e_{0},e_{1},\dots,e_{u}\}$, which represent event $\mathcal{E}$. LLMs treat this news information as a condition input and perform conditional probability predictions $P(x_{t}|x_{0:t},e_{0:u})$. Including $e_{0:u}$ provides crucial contextual information that influences the prediction of future values. This process aligns with the standard approach used by language models to interpret text, allowing for the incorporation of information about multiple events through various news contexts simultaneously to enhance prediction accuracy. In practice, we only need to integrate news text with historical time series data in a prompt engineering manner.
@@
Time series and news pre-pairing. For the initial stage of data preparation, news is retrieved to align with time series data based on matching time frequencies, horizons, and geographical areas. This synchronization ensures that insights from textual information are timely and regionally relevant. For example, to understand state-level electricity demand in Australia from 2019 to 2021, we gather local news from various Australian states and international news occurring in the same period that might directly or indirectly affect demand. In this way, potentially relevant candidate information can be roughly selected first, and such filtering can be easily completed through crawler means.
@@
Reasoning agent for news selection. We employ an LLM-based reasoning agent capable of sophisticated tasks such as conversation, reasoning, and semi-autonomous action. This agent is programmed using detailed prompts that define roles, instructions, permissions, and context, enabling it to interpret human commands and perform complex tasks. This approach condenses extensive news datasets into a refined selection of pertinent articles. It leverages its reasoning capabilities to effectively screen, categorize, and interpret news texts. We employ few-shot prompting [4] and the CoT [57] method to develop an agent that understands the context of news for forecasting. This technique enhances the model’s ability to handle multi-step, commonsense reasoning tasks essential for accurate forecasting. The agent uses multi-step prompts to break down complex problems into simpler, manageable parts. Our three-phase prompting method, illustrated in Figure 3 Part I, includes:
  1. 1.
    We develop an understanding of time series influencers, sorting them by impact (positive/negative) and duration (short/long-term), considering economic, policy, seasonal, and technological factors;
  2. 2.
    We direct the agent to filter and categorize news based on either automatically generated logic or a given reasoning logic, focusing on relevance to time series and classifying the impact (e.g., long-term and short-term) along with the rationale;
  3. 3.
    We specify the output format for the agent to organize the selected news into JSON, detailing aspects like summary, affected region, reporting time, and rationale. More details about our prompting method can be found in Appendix A.6.
The LLM agent can automatically develop an understanding of time series influencers, with the option to provide predefined reasoning logic in our models. In the automated process, the agent forms its logic through prompts designed to guide it in determining how different types of news impact a specific domain. For instance, we use open-ended questions to allow the agent to independently summarize and create its own filtering logic. User knowledge can also be incorporated into these prompts as additional reasoning, enabling the agent to generate more comprehensive logic. The agent then filters news based on the generated logic, whether fully automated or with user-provided input.
@@
Evaluation agent for reasoning updating. We also design an evaluation agent to assess and improve the effectiveness of the aforementioned news filtering. Relying solely on the reasoning agent for news selection is not optimal, as the interaction between news and time series is complex. The reasoning agent can only analyze the potential impact of different news from the perspective of news content, without knowing whether the trained time series forecasting model based on them can make accurate forecasts. The evaluation agent is deployed after the time series prediction model has been trained. The evaluation agent extends beyond simple numerical assessments of prediction accuracy by incorporating human-like logical reasoning to refine the news selection logic chain. We focus our evaluation on identifying inaccuracies potentially caused by missing news, such as unusual events or illogical reports. It observes the model’s predictive outcomes to determine if any crucial news has been overlooked and adjusts the news filtering strategy for the training data based on these results.
@@
For the evaluation agent, we structured the prompt design into three phases, as illustrated in Figure 3 Part II. In the first phase, we input the forecasting task type, the time horizon, and background information, which the agent uses to generate the steps for evaluating the prediction outcomes. In the second phase, we provide the ground truth, the discrepancies between predicted and actual series, as well as selected and historical news. The agent analyzes these inputs to identify overlooked news based on the distribution of prediction errors over time. In the third phase, the agent generates updated logic based on its analysis, guiding future news selection. After processing all validation set predictions, the reasoning agent consolidates the updated logic into a cohesive final strategy.
@@ 3.3 Overall Pipeline
We integrate the news reasoning and evaluation agents with the fine-tuning of the LLM forecasting model to enhance the quality of training data, as illustrated in Figure 4. In the first iteration, we use the LLM agent to establish news selection logic based on the domain and timing of the time series task. This logic directs the reasoning agent to filter relevant news, align it with time series data, and input it into the model for initial fine-tuning. After validating the model’s predictions with a validation set, which is randomly extracted from the training data for each iteration, the evaluation agent checks for any missing news that may have influenced the prediction. This feedback helps the reasoning agent refine the filtering logic in subsequent iterations. The cycle continues until the final iteration, where the reasoning agent consolidates all updates to create the definitive news filter for training the final model. We use the GPT-4 Turbo model as the LLM for the agents described above.